Annapolis Company has two service departments (Computer Operations & Maintenance Services). Annapolis has two production departments (Mixing Department & Packaging Department.) Annapolis uses a direct allocation method where service departments are allocated only to production departments. All allocations are based on total employees. Computer Operations has costs of $175,000 and Maintenance Services has costs of $195,000 before any allocations. What amount of Maintenance Services total cost is allocated to the Packaging Department? (round to closest whole dollar) Employees are: Computer Operations 4 Maintenance Services 3 Mixing Department 4 Packaging Department 4
Answer #1

Calculate amount allocated :

Total employees for allocation = 4+4 = 8

Packing department employee = 4

Maintenance service allocated to packing department = 195000*4/8 = 97500